OpenStudy (mackenzie2013):

The reciprocal of 5 3/7 is 5 7/3.

OpenStudy (mackenzie2013):

True or false

OpenStudy (skullpatrol):

What do you think?

OpenStudy (skullpatrol):

$$\Huge5\frac{3}{7}=5+\frac{3}{7}$$

OpenStudy (skullpatrol):

True or false?

OpenStudy (shamil98):

You can convert them into improper fractions and then compare, or do what skull has shown.

OpenStudy (mackenzie2013):

true?

OpenStudy (skullpatrol):

Yes, it is true. Now add them to get an improper fraction please.

OpenStudy (shamil98):

would it not be false tho? 5 3/7 = 38/3 the reciprocal is 3/38 5 7/3 = 22/3 22/3 does not equal 3/38

OpenStudy (skullpatrol):

Yes @shamil98 the answer to the question is false, but I haven't got there yet :)